# Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases

### 30 Jan 2024 v2

@@ C.1 Question Generation
Training a question generation (QG) model is the second step of GAIN. Because logical forms in the synthetic dataset do not have corresponding human-labeled natural language questions, evaluating QG on the synthetic dataset is difficult. Existing KBQA datasets contain questions written by crowdsourced workers for logical forms or triples, so we evaluate the question generator by generating questions on these datasets and calculate the BLEU-4 Papineni et al. (2002) and ROUGE-L Lin and Och (2004) scores [^12] Calculated by Hugging Face Evaluate. (Table 13).
@@ C.2 Retrieval Performance
The performance of schema retrieval on the GrailQA valid set and GraphQuestions test set is shown in Table 11 and 12, respectively. GAIN significantly improves the performance of schema retrieval on both GrailQA and GraphQuestions.

### 09 Feb 2024 v3

@@ C.1 Question Generation
Training a question generation (QG) model is the second step of GAIN. Because logical forms in the synthetic dataset do not have corresponding human-labeled natural language questions, evaluating QG on the synthetic dataset is difficult. Existing KBQA datasets contain questions written by crowdsourced workers for logical forms or triples, so we evaluate the question generator by generating questions on these datasets and calculate the BLEU-4 Papineni et al. (2002) and ROUGE-L Lin and Och (2004) scores (Table 13). [^12] Calculated by Hugging Face Evaluate.
@@ C.2 Retrieval Performance
The performance of schema retrieval on the GrailQA valid set and GraphQuestions test set is shown in Table 9 and 10, respectively. GAIN improves the performance of schema retrieval on both GrailQA and GraphQuestions. In particular, GAIN improves the relation recall@10 by 7.2 points on GraphQuestions.
@@ C.4 Error Analysis
Appendix D Details of Scientific Artifacts
All datasets we use are publicly available. GrailQA [^13] https://dki-lab.github.io/GrailQA Gu et al. (2021) uses CC BY-SA 4.0 license, and GraphQuestions [^14] https://github.com/ysu1989/GraphQuestions Su et al. (2016) uses CC BY 4.0 license. WebQSP [^15] https://www.microsoft.com/en-us/download/details.aspx?id=52763 Yih et al. (2016) and SimpleQuestions-Balance [^16] https://github.com/wudapeng268/KBQA-Adapter/tree/master/Data/SQB Wu et al. (2019) are also downloaded from their official release channels. We have complied with their distribution rules. These datasets involve manual construction rather than fully automated construction, which includes the review process. They contain questions about some famous individual people, but the corresponding content is available on Freebase, which is a publicly released, human-reviewed knowledge base.
Although the training data for LLMs could be quite large, the test set annotation of GrailQA is not publicly available. In addition, our experimental results on the test set and the validation set show the same trend, so the impact of the data contamination problem on the experiments of this paper could be ignored.